The osmotic pressure of human blood is 7.6 atm at 37 °C. What mass of glucose, C6H12O6, is required to make 1.00 L of aqueous solution for intravenous feeding if the solution must have the same osmotic pressure as blood at body temperature, 37 °C?


1
Expert's answer
2022-01-17T11:20:01-0500


"\Pi=iCRT"


"\Pi=7.6atm,i=1,R=0.08210L\>atm/(mol\>K)"


"T=273+37=310K"


Molar concentration (C)"=\frac{7.6}{0.08210×310}"


"=0.29898M"


Molar mass of Glucose "=180g"


Mass of glucose "=180\u00d70.29898"

"=53.8164g"